# Learning to Match Unpaired Data with Minimum Entropy Coupling

### v1

@@ Results.
ddmec(guidance=9) 62.03 $\pm$ 1.18 0.360 $\pm$ 0.0020
ddmec(guidance=7) 60.67 $\pm$ 1.01 0.353 $\pm$ 0.0035
@@ 5 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation Details
A.1 Diffusion Models Training with Reinforcement Learning
@@
A.2 Technical Details and Hyperparameters
@@
We use the SNARE-SEQ dataset available in the official code repository of (Demetci et al., 2022) https://github.com/rsinghlab/SCOT, along with the provided preprocessing steps and evaluation procedure. We find it beneficial to normalize the data by subtracting the mean and standardizing it to unit variance. In this experiment, we employ a simple MLP network as described in Table 3. The model is initially trained unconditionally for 2000 steps using Denoising Diffusion Probabilistic Models (DDPM) with $T=50$. Subsequently, we train the two conditional models for 350 steps, where each step corresponds to a single iteration of Algorithm 1, involving one policy update and one application of the denoising loss. For importance sampling, we set the clipping hyperparameter to 0.01. We use a batch size of 64, a learning rate of $5\times 10^{-3}$, and the Adam optimizer (Kingma, 2014).
Layer Details
Input Embedding Linear (dim, 128)
Condition Embedding Linear (input_dim, 128)
Time Embedding Positional Encoding (128)
Fully Connected 1 Linear (256) + ReLU
Fully Connected 2 Linear (256) + ReLU
Fully Connected 3 Linear (256) + ReLU
Fully Connected 4 Linear (256) + ReLU
Output Layer Linear (input_dim)
Table 3: Architecture of the denoising network used in Section 4.1.
@@
AFHQ
Batch Size 16
Learning Rate: $52e-$ 5
Optimizer ADAM
@@
Noise Scheduler Linear
Number of Diffusion Timesteps ($T$) 1000
Sampler DDIM
Guidance Scale (training) 7.0
Sampling steps 50
Exponential moving average Yes
@@
Reward (Monte Carlo steps) 3
Policy Update Steps 4
Importance Sampling Clipping $41e-$ 4
$\lambda_{1},\lambda_{2}$ $31e-$ 3
Gradient Accumulation 12
Gradient Clipping 1.0

### v2

@@ 50 Sampling steps
50 Sampling steps
DDMEC (guidance=9) 62.03 $\pm$ 1.18 0.360 $\pm$ 0.002
DDMEC (guidance=7) 60.67 $\pm$ 1.01 0.353 $\pm$ 0.004
100 Sampling steps
[…]
DDMEC (guidance=7) 59.22 $\pm$ 0.35 0.346 $\pm$ 0.001
[…]
50 Sampling steps
DDMEC (guidance=2.5) 40.73 $\pm$0.61 0.593 $\pm$ 0.003
[…]
100 Sampling steps
DDMEC (guidance=2.5) 38.93 $\pm$ 0.37 0.588 $\pm$ 0.002
DDMEC (guidance=2) 34.86 $\pm$ 0.70 0.549 $\pm$ 0.002
This outcome aligns with expectations, as DDMEC  is designed to reduce uncertainty and enforce adherence to marginal constraints. In Figures 3 and 2, qualitative results further confirm the performance of DDMEC  in this task. Additional results are available in Figures 6 and 9.
We investigate the effect of the guidance scale which acts as a temperature like parameter that controls conditioning strength—in our diffusion-based unpaired image translation framework. A test-time ablation study (Figure 8(a)) reveals a trade-off: higher guidance improves SSIM (structural similarity) but degrades FID (image realism), while lower guidance yields the opposite. This illustrates how guidance balances modality fidelity and mutual information maximization under the MEC framework.
@@ 5 Conclusion
Acknowledgment
Pietro Michiardi was partially funded by project MUSE-COM2 - AI-enabled MUltimodal SEmantic COMmunications and COMputing, in the Machine Learning-based Communication Systems, towards Wireless AI (WAI), Call 2022, ChistERA.
@@ Acknowledgment
This paper aims to advance Machine Learning by improving unpaired data translation, with potential benefits for applications such as multi-omics alignment in biology and image translation. However, these advancements may also increase risks associated with generative models, including the generation of deepfakes. Responsible and ethical use is advised.
Appendix A Additional Details
Figure 4: Overview of the DDMEC methodology. The phase (1) corresponds to the procedure described in Algorithm 2, Lines 2 and 3, and involves generating samples (depicted in red and blue) conditioned on inputs $x$ and $y$ drawn from their respective marginals. These samples are then used to evaluate the loss defined in Section 3.1, which is practically optimized using PPO (Fan et al., 2023). The phase (2) corresponds to Line 4 of Algorithm 2, wherein the joint consistency constraint is enforced by updating the model with the previously generated sample pairs. Both phases require coordination between the two models, alternating their roles as outlined in Algorithm 1.
A.1 Details on Swapping the Roles of the Conditional Models
In this part we provide more details about the justification of :
$\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\theta}_{X|Y}]\approx\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\phi}_{Y|Y}]$ (12)
whenever $p^{\theta}_{X,Y}=p^{\phi}_{X,Y}$.
Equation 12 reads:
$\nabla_{\theta}\int p_{Y}(y)\,p^{\theta}_{X|Y}(x|y)\log p^{\theta}_{X|Y}(x|y)\,dx\,dy.$ (13)
Moving the gradient $\nabla_{\theta}$ inside the integral and applying the chain rule, we obtain:
$\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\log p^{\theta}_{X|Y}(x|y)\,dx\,dy$
$\quad+\int p_{Y}(y)\,p^{\theta}_{X|Y}(x|y)\,\nabla_{\theta}\left(\log p^{\theta}_{X|Y}(x|y)\right)\,dx\,dy.$ (14)
The second term simplifies to zero:
$\int p_{Y}(y)\,p^{\theta}_{X|Y}(x|y)\,\nabla_{\theta}\left(\log p^{\theta}_{X|Y}(x|y)\right)\,dx\,dy$ $=\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\,dx\,dy$
$=\nabla_{\theta}\int p_{Y}(y)\,p^{\theta}_{X|Y}(x|y)\,dx\,dy$
$=\nabla_{\theta}1=0.$ (15)
Assuming $p^{\theta}_{X,Y}=p^{\phi}_{X,Y}$, $p^{\theta}_{X}=p_{X}$, and $p^{\phi}_{Y}=p_{Y}$, the first term rewrites as:
$\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\log\frac{p^{\phi}_{Y|X}(y|x)p_{X}(x)}{p_{Y}(y)}\,dx\,dy$
$=\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\left(\log p^{\phi}_{Y|X}(y|x)+\log p_{X}(x)-\log p_{Y}(y)\right)\,dx\,dy$
$=\nabla_{\theta}\int p_{Y}(y)\,p^{\theta}_{X|Y}(x|y)\log p^{\phi}_{Y|X}(y|x)\,dx\,dy,$ (16)
which corresponds to the right-hand side (r.h.s) of Equation 12.
Indeed, the additional terms vanish:
$\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\log p_{X}(x)\,dx\,dy$ $=\int\nabla_{\theta}\left(p^{\theta}_{X}(x)\right)\log p_{X}(x)\,dx$
$=\int\nabla_{\theta}\left(p_{X}(x)\right)\log p_{X}(x)\,dx=0,$ (17)
and similarly,
$\int p_{Y}(y)\,\nabla_{\theta}\left(p^{\theta}_{X|Y}(x|y)\right)\log p_{Y}(y)\,dx\,dy=0.$ (18)
A.2 Diffusion Models Training with Reinforcement Learning
@@
A.3 Technical Details and Hyperparameters
The source  [^1] https://github.com/MustaphaBounoua/ddmec is publicly available.
@@ A.3 Technical Details and Hyperparameters
For the PBMC and BM datasets, we utilize the preprocessed versions provided in the official code repository of Singh et al. (2023) [^2] https://github.com/AkashCiel/scTopoGAN, along with the accompanying evaluation protocols. Prior to training, we normalize the data by subtracting the mean and scaling to unit variance, while applying outlier mitigation. The model is first trained unconditionally using a DDPM for 100,000 steps with $T=1000$ diffusion steps. To stabilize training, we additionally incorporate a nearest neighbor retrieval step after each generation, where generated samples are projected back to the closest real data points using Euclidean distance. Subsequently, we train the two conditional models, following the procedure outlined in Algorithm 1. For each training step, we use a batch size of 256 and perform four gradient updates corresponding to line 3, followed by four updates for Line 4. We find it beneficial to accumulate generated samples during training and reuse them in optimizing Line 4. We use a simple MLP network with skip connections and use the Adam optimizer (Kingma, 2014) with a learning rate of $1\times 10^{-4}$. The KL divergence regularization weight is set to $\lambda=0.01$ for PBMC and $\lambda=0.02$ for BM.
@@
AFHQ CelebA-HQ
Batch Size 16 16
Learning Rate: 2e52e-5 2e52e-5
Optimizer ADAM ADAM
@@
Noise Scheduler Linear Linear
Number of Diffusion Timesteps ($T$) 1000 1000
Sampler DDIM DDIM
Guidance Scale (training) 7.0 7.0
Sampling steps 50 50
Exponential moving average Yes Yes
@@
Reward (Monte Carlo steps) 3 3
Policy Update Steps 4 4
Importance Sampling Clipping 1e41e-4 1e41e-4
$\lambda_{1},\lambda_{2}$ 1e31e-3 4e34e-3
Gradient Accumulation 12 12
Gradient Clipping 1.0 1.0
@@ Appendix B Additional Results
B.1 SNARE-seq additional experiments
In this experiment we use the SnaReSeq (Chen et al., 2019) dataset, which links chromatin accessibility with gene expression data on a mixture of four cells types. We use the same preprocessing procedures detailed in (Demetci et al., 2022), which deal with filtering spurious data affected by technical errors, and normalization. Data samples have 1–1 correspondence information, which constitute the groud-truth information used for our performance evaluation. We use the average “fraction of samples closer than the true match (foscttm)” metric introduced by Liu et al. (2019): given a sample in one domain, this amounts to compute the fraction of samples that are positioned more closely to it than its true match after pairing. Results report the average foscttm across all samples, where lower values indicate better performance. We also report the label transfer accuracy as done by Cao et al. (2020), which measures how well sample labels are transferred between domains based on neighborhood alignment. A $k$-nearest neighbor classifier is trained on one domain and used to predict labels in the other. In this experiment, DDMEC  is trained once, and inference is conducted five times with different seeds. the foscttm metric, and is on-par with the best method in terms of accuracy. In this experiment, DDMEC  is trained once, and inference is conducted five times with different seeds. Unlike other methods, DDMEC  is conceptually different as it generates samples rather than learning a deterministic, 1-1 mapping. To compute the different metrics, given a sample from one modality, we use DDMEC  to generate a coupling to the other modality, then select the nearest sample from the dataset based on Euclidean distance.
SNAREseq
FOS $\downarrow$ Acc $\uparrow$
[…]
DDMEC 0.147 98.6
Table 4: Performance results on SNAREseq dataset
@@ B.1 SNARE-seq additional experiments
(a) In the guidance scale ablation study, we report the FID and SSIM as a function of the guidance scale on the CelebA-HQ dataset. We notice that an increase in the guidance scale results in more information transfer between the two modalities, leading to a worse FID, and vice versa.
Source
Output
Source
Output
[…]